### v1

# The MLE is minimax optimal for LGC

Abstract
We revisit the recently introduced Local Glivenko-Cantelli setting, which studies distribution-dependent uniform convegence rates of the Maximum Likelihood Estimator (MLE). In this work, we investigate generalizations of this setting where arbitrary estimators are allowed rather than just the MLE. Can a strictly larger class of measures be learned? Can better risk decay rates be obtained? We provide exhaustive answers to these questions — which are both negative, provided the learner is barred from exploiting some infinite-dimensional pathologies. On the other hand, allowing such exploits does lead to a strictly larger class of learnable measures.
October 2, 2024
1 Introduction
Cohen and Kontorovich (2023) initiated the study of the local Glivenko-Cantelli setting: laws of large numbers that are uniform over a function class but rather than being universal over all distributions, feature a delicate dependence of the risk decay on the (local) sampling measure. This naturally led to the binomial empirical process: for a fixed $p\in[0,1]^{\mathbb{N}}$ and each $n\in\mathbb{N}$, we have a sequence of independent $Y_{j}\sim\operatorname{Binomial}(n,p_{j})$, which are centered and normalized to obtain $\bar{Y}_{j}:=n^{-1}Y_{j}-p_{j}$. The object of interest is the expected uniform absolute deviation:
$\Delta_{n}$ ::=:= $\mathop{\mathbb{E}}\sup_{j\in\mathbb{N}}|\bar{Y}_{j}|.$ (1)
More generally, one could imagine fixing a distribution $\mu$ on $\left\{0,1\right\}^{\mathbb{N}}$, sampling $X^{(1)},X^{(2)},\ldots,X^{(n)}$ i.i.d. from $\mu$, and estimating $p:=\mathop{\mathbb{E}}X^{(1)}$ via the Maximum-Likelihood Estimator (MLE) $\hat{p}:=\frac{1}{n}\sum_{i=1}^{n}X^{(i)}$. In the case where $\mu$ is a product measure (that is, the components of the vector $X\sim\mu$ are mutually independent), $\mathop{\mathbb{E}}\left\|\hat{p}-p\right\|_{\infty}$ recovers the expression in (1). Despite its austere appearance, the binomial empirical process with independent coordinates $Y_{j}$ under $\ell_{\infty}$-norm deviation already captures much of the richness of problem. Extensions to more general product distributions $\mu$ over $[0,1]^{\mathbb{N}}$ are straightforward (Blanchard and Voráček, 2024, Corollary 6) and the behavior under $\ell_{r}$ norms for $r<\infty$ is considerably simpler (Proposition 7 ibid.). Finally, the in-expectation bounds are readily converted to high-probability tail bounds (Proposition 9 ibid.), and all of the upper bounds stated for product measures hold verbatim for arbitrary correlations.
For the purpose of analyzing (1), Cohen and Kontorovich showed that there is no loss of generality in restricting $p$ to the set $[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}_{\downarrow 0}$, consisting of all $p\in[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}$ with $p_{j}\downarrow 0$. They defined $\mathsf{LGC}\subset[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}_{\downarrow 0}$ as the family of $p$ for which $\Delta_{n}\mathop{\longrightarrow}\limits_{n\to\infty}0$ and showed that $\mathsf{LGC}$ consists of exactly those $p$ for which
$T(p)$ ::=:= $\sup_{j\in\mathbb{N}}\frac{\log(j+1)}{\log(1/p_{j})},\qquad p\in[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}_{\downarrow 0}$ (2)
is finite. They also characterized up to constants the asymptotic decay of $\Delta_{n}$ (whenever $T(p)<\infty$) via the functional
$S(p)$ ::=:= $\sup_{j\in\mathbb{N}}p_{j}\log(j+1),\qquad p\in[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}_{\downarrow 0},$ (3)
establishing that $\Delta_{n}(p)$ decays as $\sqrt{S(p)/n}$. Additional finite-sample bounds provided therein were tightened by Blanchard and Voráček (2024) as follows:
$\Delta_{n}(p)$ $\asymp 1\land\left(\sqrt{\frac{S(p)}{n}}+\sup_{j\geq 1}\frac{\log(j+1)}{n\log\left(2+\frac{\log(j+1)}{np_{j}}\right)}\right),$ if $\text{if }n\cdot\sup_{j\geq 1}2jp_{j}>1,$
$\Delta_{n}(p)$ $\asymp\frac{1}{n}\land\sum_{j\geq 1}p_{j},$ otherwise.
In a later work, Blanchard et al. (2024) extended some of the analysis to the much more difficult case where $\mu$ is not a product measure (i.e., the coordinates of $X\sim\mu$ have correlations). In the present paper, we return to the product-measure case and investigate a different extension: How does $\Delta_{n}$ behave if rather than restricting the estimator to the MLE $\hat{p}$, we allow arbitrary estimators $\tilde{p}$?
Formally, an estimator $\tilde{p}$ is any mapping from ($(\left\{0,1\right\}^{\mathbb{N}})^{n}$ to $[0,1]^{\mathbb{N}}$. Any $p\in[0,1]^{\mathbb{N}}$ induces the product measure
$\mu=\mu(p)=\operatorname{Bernoulli}(p_{1})\otimes\operatorname{Bernoulli}(p_{2})\otimes\ldots$ (4)
on $\left\{0,1\right\}^{\mathbb{N}}$. If $X^{(1)},X^{(2)},\ldots,X^{(n)}$ are sampled i.i.d. from $\mu$, then these induce $\tilde{\Delta}_{n}:=\mathop{\mathbb{E}}\left\|\tilde{p}-p\right\|_{\infty}$. We say that a family of product distributions induced by $\mathcal{P}\subset[0,1]^{\mathbb{N}}$ is learnable by $\tilde{p}$ if $\tilde{\Delta}_{n}\mathop{\longrightarrow}\limits_{n\to\infty}0$ for each $p\in\mathcal{P}$, and just learnable if it is learnable by some $\tilde{p}$. (Since the sequence $p$ fully determines the measure $\mu(p)$, it is fitting to say that $\tilde{p}$ “learns” $p$ — and hence also $\mu(p)$.)
This general setting immediately raises the natural questions: Can $\mathsf{LGC}$ be expanded to a larger learnable family via some estimator $\tilde{p}$ different from the MLE? Can some estimator $\tilde{p}$ achieve better decay rates for $\tilde{\Delta}_{n}$ than the MLE?
Our contributions.
Modulo some technical caveats, we resolve both questions above in the negative. If the learner is barred from exploiting some pathological quirks of the infinite-dimensional setting, then essentially $\mathsf{LGC}$ as defined above is the largest learnable family (Theorem 1). Furthermore, the MLE achieves the minimax risk decay rate over non-pathological distribution families (Theorem 2). Finally, in Theorem 3 we show that non-trivial extensions of $\mathsf{LGC}$ become possible once the restrictions are relaxed.
Related work.
Estimating the mean of a high-dimensional distribution from independent draws is among the most basic problems of statistics. Much of the earlier theory has focused on obtaining efficient estimators $\hat{m}_{n}$ of the true mean $m$ and analyzing the decay of $\left\|\hat{m}_{n}-m\right\|_{2}$ as a function of sample size $n$, dimension $d$, and various moment assumptions on $X$ (Catoni, 2012; Devroye et al., 2016; Lugosi and Mendelson, 2019a, b; Cherapanamjeri et al., 2019, 2020; Diakonikolas et al., 2020; Hopkins, 2020; Lugosi and Mendelson, 2021; Lee and Valiant, 2022). For $d$-dimensional distributions $\mu$ on $\{0,1\}^{d}$, Chernoff and union bounds yield $\Delta_{n}(\mu)\lesssim\sqrt{\ln(d+1)/n}$ for the MLE, and a simple information-theoretic argument shows that this is minimax-optimal up to constants (Cohen and Kontorovich, 2023, Proposition 1). Cohen and Kontorovich further motivated their choice of the $\ell_{\infty}$ norm as the most interesting of all the $\ell_{r}$ norms, in a well-defined sense (see Blanchard and Voráček (2024, Proposition 7)). Blanchard and Voráček (2024) fully closed the gaps in the analysis of Cohen and Kontorovich, and Blanchard et al. (2024) took the first nontrivial steps in analyzing non-product sampling distributions.
Notation.
The measure-theoretic subtleties of defining distributions on $\left\{0,1\right\}^{\mathbb{N}}$ are addressed in Cohen and Kontorovich (2023). Our logarithms will always be base $\mathrm{e}$ by default; other bases will be explicitly specified. The natural numbers are denoted by =$\mathbb{N}=\left\{1,2,3,\dots\right\}$ and for $k\in\mathbb{N}$, we write $[k]=\left\{i\in\mathbb{N}:i\leq k\right\}$. The floor and ceiling functions, ${\left\lfloor t\right\rfloor}$, ${\left\lceil t\right\rceil}$, map $t\in\mathbb{R}$ to its closest integers below and above, respectively; also, $s\vee t:=\max\left\{s,t\right\}$, $s\wedge t:=\min\left\{s,t\right\}$. Unspecified constants such as $c,c^{\prime}$ may change value from line to line. We use superscripts to denote distinct random vectors and subscripts to denote indices within a given vector. Thus, if $X^{(1)},,\ldots,X^{(n)}$ are independent copies of $X$, then $X^{(i)}_{j}$ denotes the $j$th entry of the $i$th copy.
When considering the MLE as the sole estimator (as in previous works), no generality was lost in restricting the range of $p$ to $[0,\frac{1}{2}]$ and assuming sequences monotonically decreasing to 00 (i.e., $[0,{\textstyle\frac{1}{2}}]^{\mathbb{N}}_{\downarrow 0}$). The definitions of $T$ and $S$ in (2, 3) were based on this assumption. In this work, we will need their slightly generalized versions. With the convention $\dot{x}:=\min\left\{x,1-x\right\}$, we define
$T(p)$ ::=:= $\inf_{\sigma:\mathbb{N}\to\mathbb{N}}\sup_{j\in\mathbb{N}}\frac{\log(j+1)}{\log(1/\dot{p}_{\sigma(j)})},$ (5)
$S(p)$ ::=:= $\inf_{\sigma:\mathbb{N}\to\mathbb{N}}\sup_{j\in\mathbb{N}}\dot{p}_{\sigma(j)}\log(j+1),$ (6)
for $p\in[0,1]^{\mathbb{N}}$, where the infimum is over all permutations $\sigma$ over $\mathbb{N}$. Whenever $\dot{p}_{j}\to 0$, a unique non-increasing permutation ${\dot{p}}^{\downarrow}$ exists, and it is easily seen to be the one achieving both infima above; in this case, the definitions in (5, 6) coincide with those in (2, 3).
Any $p\in[0,1]^{\mathbb{N}}$ defines the product measure $\mu=\mu(p)$ as in (4). An estimator $\tilde{p}$ and its induced deviation $\tilde{\Delta}_{n}$ are defined just above (4), and the learnability of a family $\mathcal{P}\subset[0,1]^{n}$ is defined just below it.
We say that a family $\mathcal{P}\subset[0,1]^{\mathbb{N}}$ is decaying if ${\displaystyle\lim_{j\to\infty}\dot{p}_{j}}=0$ for all $p\in\mathcal{P}$. For $p\in[0,1]^{\mathbb{N}}$ and $b\in\left\{-1,1\right\}^{\mathbb{N}}$, we say that $p^{\prime}=p^{\prime}(p,b)\in[0,1]^{\mathbb{N}}$ is a $b$-reflection of $p$ about $\frac{1}{2}$ if $p^{\prime}_{j}=b_{j}(p_{j}-\frac{1}{2})+\frac{1}{2}$ for all $j\in\mathbb{N}$. We say that $\mathcal{P}\subset[0,1]^{\mathbb{N}}$ is strongly symmetric about $\frac{1}{2}$ if $p\in\mathcal{P}\implies p^{\prime}(p,b)\in\mathcal{P}$ holds for all $p\in\mathcal{P}$ and $b\in\left\{-1,1\right\}^{\mathbb{N}}$.
The family $\mathsf{LGC}\subset[0,\frac{1}{2}]^{\mathbb{N}}$ was defined in Cohen and Kontorovich (2023) as the one learnable by the MLE $\hat{p}$, and characterized therein as consisting precisely of those $p\in[0,\frac{1}{2}]^{\mathbb{N}}$ for which $T(p)<\infty$. Since in this work we do not restrict the range of $p$ to $[0,\frac{1}{2}]$, we define
$\dot{\mathsf{LGC}}$ ::=:= $\left\{p\in[0,1]^{\mathbb{N}}:T(p)<\infty\right\},$ (7)
for $T$ as defined in (5). It is straightforward to extend the arguments of Cohen and Kontorovich (2023) to show that $\dot{\mathsf{LGC}}$ consists precisely of those $p\in[0,1]^{\mathbb{N}}$ for which the MLE $\hat{p}$ yields $\Delta_{n}\to 0$.
2 Main Results
Our first result may be informally summarized thus: morally speaking, $\mathsf{LGC}$ is the largest family that is learnable by any fixed estimator.
Theorem 1 (expanding $\mathsf{LGC}$).
Suppose that $\mathcal{P}\subset[0,1]^{\mathbb{N}}$ defines a family of product distributions as in (4) and furthermore
  1. 1.
    $\mathcal{P}$ is decaying
  2. 2.
    $\mathcal{P}$ is strongly symmetric about $\frac{1}{2}$
  3. 3.
    $\mathcal{P}$ is learnable.
Then $\mathcal{P}\subseteq\dot{\mathsf{LGC}}$.
Remark.
Strong symmetry about $\frac{1}{2}$ forces the sequences in $\mathcal{P}$ to be “generic” and prevents the learner from beating the MLE by exploiting some special structure. Note that this condition is very much absent in Theorem 3, where indeed such exploits become possible.
Having established that (modulo pathologies) $\mathsf{LGC}$ is the largest learnable family, we next show that the MLE is nearly minimax-optimal for this family.
Theorem 2 (Minimax bound).
There exist universal constants $c,c^{\prime},C>0$ such that the following holds. For $n\in\mathbb{N}$ and $s,t>0$ satisfying $\frac{c^{\prime}\log n}{n}\leq\frac{s}{t}\leq\mathrm{e}^{-1}$, let
$\mathcal{P}_{s,t}:=\left\{p\in[0,1]^{\mathbb{N}}:S(p)\leq s\,\wedge\,T(p)\leq t\right\}.$
Then, whenever $\frac{c^{\prime}\log n}{n}\leq\frac{s}{t}\leq\mathrm{e}^{-1}$ and
$n\geq\frac{\frac{t^{2}}{Cs}\log\frac{t}{s}}{\frac{t}{s}\log\frac{t}{s}\cdot\mathrm{e}^{-\frac{t\log\frac{t}{s}}{\log 2}}-1},$
we have
$\inf_{\tilde{p}}\sup_{p\in\mathcal{P}_{s,t}}\mathop{\mathbb{E}}\sup_{j\in\mathbb{N}}|\tilde{p}_{j}-p_{j}|\geq 1\wedge\left(c\sqrt{\frac{s}{n}}\,\vee\,Q(t,s)\cdot\frac{t}{n}\right),$ (8)
where the infimum is over all estimators $\tilde{p}$ that are based on $n$ i.i.d. samples drawn from $p$, and
$Q(t,s)=C\left(1+\frac{\log\frac{t}{s}}{\log\log\frac{t}{s}}\right)^{-1}.$
Remark.
The logarithmic factor and restrictions on the range of $n$ are likely artifacts of the argument, which we kept streamlined for space and readability. We look forward to removing both in the extended version.
Finally, we show that if the learner is allowed to “cheat” by exploiting the information contained in the infinitely many bits of each example $X^{(i)}$, then $\mathsf{LGC}$ can indeed be non-trivially expanded. Let us elaborate a bit on the nature of these exploits. The elements of $\mathsf{LGC}$ have a “generic,” unstructured flavor: knowing the values of $p_{j}$ for $j\in[N]$ reveals no useful information regarding the remaining $j>N$; all the learner knows is that these must decay as some power of $j$ in order to be in $\mathsf{LGC}$. On the other hand, one might consider adjoining a “structured” sequence to $\mathsf{LGC}$, such as $p=(\frac{1}{2},\frac{1}{2},\ldots)$. Because a single $X^{(i)}\sim\mu$ provides a bit drawn from each of the $\operatorname{Bernoulli}(p_{j})$, the learner (as we show below) is able to first test whether the unknown sequence has the given structure (in this case, whether it was generated by $p\equiv\frac{1}{2}$) and if not, then reverts to the standard MLE for learning the unstructured sequences in $\mathsf{LGC}$.
Theorem 3 (Relaxing decay and symmetry).
Define the family $\mathrm{const}\subset[0,1]^{\mathbb{N}}$ by
$\mathrm{const}$ ::=:= $\left\{(c,c,\dots):c\in[0,1]\right\}.$
Then $\mathcal{P}=\dot{\mathsf{LGC}}\cup\mathrm{const}$ is learnable, meaning that there exists an estimator $\tilde{p}$ such that $\tilde{\Delta}_{n}(p)\to 0$ for all $p\in\mathcal{P}$.
Remark.
The techniques of Theorem 3 are applicable considerably more broadly than just to the family $\mathcal{Q}=\mathrm{const}$. For example, the argument can be easily adapted to show that $\dot{\mathsf{LGC}}\cup\mathcal{Q}$ is learnable for any finite $\mathcal{Q}$.
Open problems.
Two natural directions for future study are extensions of Theorems 1 and 3. For the former, it is likely that the conditions on $\mathcal{P}$ are too stringent and can be significantly relaxed; in particular, requiring that $\mathcal{P}$ be decaying is quite probably unnecessary. Thus, we seek a larger family $\mathcal{P}^{\prime}$ whose learnability implies $\mathcal{P}^{\prime}\subseteq\dot{\mathsf{LGC}}$. Regarding Theorem 3, we again anticipate the existence of considerably richer families $\mathcal{Q}$ for which $\dot{\mathsf{LGC}}\cup\mathcal{Q}$ is learnable. One such family is proposed in the conjecture below.
Conjecture.
Let $\mathcal{Q}\subset[0,1]^{\mathbb{N}}$ be a countable family of sequences with the following property: for each $q,q^{\prime}\in\mathcal{Q}$, there is an $\varepsilon>0$ and an infinite $J\subset\mathbb{N}$ such that $|q_{j}-q^{\prime}_{j}|>\varepsilon$ for all $j\in J$. Then $\mathsf{LGC}\cup\mathcal{Q}$ is learnable.
3 Proofs
3.1 Proof of Theorem 1
Assume, for the sake of contradiction, that there exists an estimator $\tilde{p}$ and a family $\mathcal{P}\subset[0,1]^{\mathbb{N}}$ satisfying the conditions of the theorem, such that $\mathcal{P}$ is learnable by $\tilde{p}$ but there exists a $p^{*}\in\mathcal{P}\setminus\dot{\mathsf{LGC}}$. Based on $p^{*}$, we will construct a family $\mathcal{P}^{*}\subset\mathcal{P}$ and argue that $\tilde{\Delta}_{n}(p)\to 0$ cannot hold for all $p\in\mathcal{P}^{*}$.
Since $\mathcal{P}$ is decaying and strongly symmetric about $\frac{1}{2}$, for any $p\in\mathcal{P}$ and any sign vector $b\in\{-1,1\}^{\mathbb{N}}$, the $b$-reflection $p^{\prime}(p,b)$ defined by $p^{\prime}_{j}=b_{j}(p_{j}-\frac{1}{2})+\frac{1}{2}$ also belongs to $\mathcal{P}$.
Consider the following randomized experiment:
  • Let $Y=(Y_{j})_{j\in\mathbb{N}}$ be a sequence of independent Rademacher random variables, i.e., $\mathbb{P}(Y_{j}=1)=\mathbb{P}(Y_{j}=-1)=\frac{1}{2}$.
  • Define $p^{(Y)}\in[0,1]^{\mathbb{N}}$ as the $Y$-reflection of $p^{*}$ about $\frac{1}{2}$, i.e. $p_{j}^{(Y)}=Y_{j}(p_{j}^{*}-\frac{1}{2})+\frac{1}{2}$.
  • Generate $n$ independent draws $X^{(1)},\dots,X^{(n)}\in\{0,1\}^{\mathbb{N}}$ from the product distribution $\mu(p^{(Y)})$ as in (4).
The assumption that $\dot{p}_{j}^{*}\in[0,\frac{1}{4}]^{\mathbb{N}}$ incurs no loss of generality, since the decay condition implies that $\dot{p}_{j}^{*}\leq\frac{1}{4}$ will hold for all sufficiently large $j$.
We follow the standard reduction from the harder problem of estimating $p^{(Y)}$ to the easier problem of recovering the sign vector $y\in\{-1,1\}^{\mathbb{N}}$ that defines the $y$-reflection $p^{(Y)}$. By the Neyman-Pearson lemma, an optimal estimator $\hat{y}$ is one that minimizes the posterior probability of error, i.e.,
$\hat{y}=\operatorname*{arg\,min}_{y\in\{-1,1\}^{\mathbb{N}}}\mathbb{P}(Y\neq y\mid X=x),$
where $Y$ is the random sign vector and $X=\left(X^{(1)},\dots,X^{(n)}\right)$ denotes the observed data.
Now
$1-\mathbb{P}(Y\neq y\mid X=x)$ $=$ $\mathbb{P}(Y=y\mid X=x)$
$=$ $\mathbb{P}(Y_{1}=y_{1}\mid X=x)\mathbb{P}(Y_{2}=y_{2}\mid X=x,Y_{1}=y_{1})\cdot$
$\mathbb{P}(Y_{3}=y_{3}\mid X=x,Y_{1}=y_{1},Y_{2}=y_{2})\cdot\ldots.$
Since the $Y_{j}$ are mutually independent each of the factors above has the simpler form
$\mathbb{P}(Y_{k}=y_{k}\mid X=x,Y_{1}=y_{1},\ldots,Y_{k-1}=y_{k-1})$ $=$ $\mathbb{P}(Y_{k}=y_{k}\mid X=x).$
We conclude that he events $E_{j}=\left\{Y_{j}\neq y_{j}\mid X\right\}$ are mutually independent. Thus
$\mathbb{P}(Y\neq y\mid X=x)$ $=$ $\mathbb{P}\left(\bigcup_{j\in\mathbb{N}}E_{j}\right)$
$=$ $\lim_{N\to\infty}\mathbb{P}\left(\bigcup_{j=1}^{N}E_{j}\right)$
$=$ $\lim_{N\to\infty}\alpha_{N}\left(\mathbb{P}(E_{1}),\mathbb{P}(E_{2}),\ldots,\mathbb{P}(E_{N})\right),$
where the second equality holds by regularity of probability measures (Kechris, 1995, Theorem 17.10), and $\alpha_{N}:[0,1]^{N}\to[0,1]$ is the inclusion-exclusion function defined inductively by $\alpha_{1}(x)=1$ and
$\alpha_{N+1}(x_{1},x_{2},\ldots,x_{N},x_{N+1})$ $=$ $x_{N+1}+(1-x_{N+1})\alpha_{N}(x_{1},x_{2},\ldots,x_{N}).$
By Kontorovich (2012, Lemma 4.2), $\alpha_{N}$ is monotonically increasing in each argument. Hence, the optimal estimator may minimize each $\mathbb{P}(E_{j})$ individually — and so we may define $A_{j}$ as the estimator for the $j$-th coordinate, where $A_{j}:\{0,1\}^{\mathbb{N}\times n}\to[0,1]$ is any mapping from the $j$-th row of the data matrix to an estimate of $p_{j}$. Let $B_{j}$ be the event that $A_{j}$ and $p_{j}^{(Y)}$ belong to different intervals, i.e., either $A_{j}\in[0,\frac{1}{2})$ and $p_{j}^{(Y)}\in(\frac{1}{2},1]$ or vice versa. To establish a lower bound on the error of any estimator, consider the minimax risk:
$\inf_{A}\sup_{p\in\mathcal{P}}\mathop{\mathbb{E}}_{X}\sup_{j\in\mathbb{N}}|A_{j}-p_{j}|$ $\geq\inf_{A}\mathop{\mathbb{E}}_{Y}\mathop{\mathbb{E}}_{X}\sup_{j\in\mathbb{N}}|A_{j}-p^{(Y)}|$
$\geq\inf_{A}\mathop{\mathbb{E}}_{Y}\mathop{\mathbb{E}}_{X}\sup_{j\in\mathbb{N}}\mathbf{1}\{B_{j}\}\left|\frac{1}{2}-p^{(Y)}\right|$
$\geq\frac{1}{4}\inf_{A}\mathop{\mathbb{E}}_{Y}\mathop{\mathbb{E}}_{X}\sup_{j\in\mathbb{N}}\mathbf{1}\{B_{j}\}$
$=\frac{1}{4}\inf_{A}\mathbb{P}_{Y,X}\left(\bigcup_{j\in\mathbb{N}}B_{j}\right)$
[…]
By the Neyman-Pearson lemma, the optimal choice of $\hat{y}_{j}$ is according to the majority vote111 The issue of optimally breaking ties or allowing randomized decision rules is somewhat delicate and is exhaustively addressed in Kontorovich and Pinelis (2019, Eq. (2.7)). In our setting, these do not affect the probability of error. of the vector $(X^{(1)}_{j},\dots,X^{(n)}_{j})$. In the event that $(X^{(1)}_{j},\dots,X^{(n)}_{j})=(1,1,\dots,1)$, but $1-\dot{p}_{j}^{*}\neq p_{j}^{(Y)}$, the estimator makes a mistake. The probability of such event, conditioned on the other random variables $X_{j^{\prime}},Y_{j^{\prime}}$ where $j^{\prime}\neq j$, is exactly $\frac{1}{2}(\dot{p}_{j}^{*})^{n}$. Since we assumed $p^{*}\notin\mathsf{LGC}$ and thus $\dot{p}^{*}\notin\mathsf{LGC}$, we have $T(\dot{p}^{*})=\infty$. Since $\dot{p}_{j}^{*}\to 0$, we may assume without loss of generality that it is decreasing monotonically. By Cohen and Kontorovich (2023, Lemma 3), it follows that for all $n\in\mathbb{N}$, we have
$\sum_{j=1}^{\infty}(\dot{p}_{j}^{*})^{n}=\infty.$
Since the events of $\hat{y}_{j}$ being wrong are mutually independent, the second BorelÿCantelli lemma implies that almost surely at least one of them will occur. It follows that $\tilde{\Delta}_{n}(p^{(Y)})\geq\frac{1}{4}$, contradicting the learnability assumption. $\square$
3.2 Proof of Theorem 2
We follow the standard approach of reducing the minimax lower bound problem to one over a finite set of hypotheses. For $2\leq J\in\mathbb{N}$ and $20\leq q\leq q^{\prime}\leq 1/$ 2 to be chosen below, we consider $J+1$ profiles $p^{(k)}\in[0,\frac{1}{2}]^{\mathbb{N}}$ for $k\in[J+1]$. For $k=J+1$ we take the step profile
$p_{j}^{(J+1)}=\begin{cases}q,&j\in[J+1],\\ 0,&j>J+1,\end{cases}$
and for $1\leq k\leq J$ we take the same step profile but with an additional bump at position $k$,
$p_{j}^{(k)}=\begin{cases}q,&j\in[J+1]\text{ and }j\neq k,\\ q^{\prime},&j=k,\\ 0,&j>J+1.\end{cases}$
Note that for all $k\neq\ell\in[J+1]$ we have $\lVert p^{(k)}-p^{(\ell)}\rVert_{\infty}=|q^{\prime}-q|$. In addition, for $k=J+1$,
$S(p^{(J+1)})=q\log(J+1)\qquad\text{and}\qquad T(p^{(J+1)})=\frac{\log(J+1)}{\log\frac{1}{q}},$
and for $k\in[J]$,
$S(p^{(k)})$ $=\max\left\{q\log(J+1),q^{\prime}\log 2\right\}\quad\text{and}\quad T(p^{(k)})=\max\left\{\tfrac{\log(J+1)}{\log\tfrac{1}{q}},\tfrac{\log 2}{\log\tfrac{1}{q^{\prime}}}\right\}.$
Given $s\leq\frac{t}{\mathrm{e}}$ as in the Theorem statement, we choose $q\in[0,\frac{1}{2}]$ and $J$ as
$q=\frac{1}{\frac{t}{s}\log\frac{t}{s}}\qquad\text{and}\qquad\log(J+1)=t\log\frac{t}{s}.$
Then
$S(p^{(J+1)})$ $=q\log(J+1)=s,$ (9)
$[3pt]T(p^{(J+1)})$ $=\frac{\log(J+1)}{\log\frac{1}{q}}=t\cdot\left(1+\frac{\log\frac{t}{s}}{\log\log\frac{t}{s}}\right)^{-1}\leq t.$ (10)
Below we set $q^{\prime}\leq 1/2$ such that $q^{\prime}\geq q$ and for all $k\in[J]$,
$S(p^{(k)})\leq S(p^{(J+1)})=s\quad\text{and}\quad T(p^{(k)})$ $\leq T(p^{(J+1)})\leq t.$ (11)
Thus, $p^{(k)}\in\mathcal{P}_{s,t}$ for all $k\in[J+1]$ and
$\inf_{\tilde{p}}\sup_{p\in\mathcal{P}_{s,t}}\mathop{\mathbb{E}}\sup_{j\in\mathbb{N}}|\tilde{p}_{j}-p_{j}|\geq\inf_{\tilde{p}}\max_{k\in[J+1]}\mathop{\mathbb{E}}_{X_{n}\sim\mu^{(k,n)}}\lVert\tilde{p}(X_{n})-p^{(k)}\rVert_{\infty}.$ (12)
To lower bound the right-hand side of (12) we apply the generalized Fano method. For $k\in[J+1]$, let $\mu^{(k)}=\mu(p^{(k)})$ be the product measure over $\left\{0,1\right\}^{\mathbb{N}}$ as defined in (4) and note that $\mathop{\mathbb{E}}_{X\sim\mu^{(k)}}\{X\}=p^{(k)}$. We denote by $\mu^{(k,n)}$ the product measure of $n$ independent copies of $X\sim\mu^{(k)}$. We invoke Lemma 1 with the $J+1$ measures $(\nu_{1},\dots,\nu_{J+1})=(\mu^{(1,n)},\ldots,\mu^{(J+1,n)})$, the distance function $\rho=\lVert\cdot\rVert_{\infty}$, and the parameters $\theta(\mu^{(k,n)})=\mathop{\mathbb{E}}_{X\sim\mu^{(k)}}\{X\}=p^{(k)}$ for $k\in[J+1]$. Note that $\rho(\theta(\mu^{(k,n)}),\theta(\mu_{\ell}^{n}))=|q^{\prime}-q|$ for all $k\neq\ell\in[J+1]$ and that
$D_{\textup{KL}}(\mu^{(k,n)}\lVert\mu_{\ell}^{n})$ $\leq n(h(q\lVert q^{\prime})+h(q^{\prime}\lVert q)),$
where
$h(q\lVert q^{\prime})=q\log\frac{q}{q^{\prime}}+(1-q)\log\frac{1-q}{1-q^{\prime}}.$
Then Lemma 1 implies
$\inf_{\tilde{p}}\max_{k\in[J+1]}\mathop{\mathbb{E}}_{X_{n}\sim\mu^{(k,n)}}\lVert\tilde{p}(X_{n})-p^{(k)}\rVert_{\infty}$ $\geq\frac{q^{\prime}-q}{2}\left(1-\left(\frac{n(h(q\|q^{\prime})+h(q^{\prime}\|q))+\log 2}{\log(J+1)}\right)\right).$ (13)
We now fix $q^{\prime}(q)=q^{\prime}(q,n,J)\geq q$ to be the solution to the equation
$h(q\|q^{\prime}(q))+h(q^{\prime}(q)\|q)=\frac{\log(J+1)}{2Cn}.$ (14)
Below we verify that (11) indeed holds with this choice of $q^{\prime}(q)$. Substituting (14) into (13), we obtain the lower bound
$\frac{q^{\prime}(q)-q}{2}\left(1-\left(\frac{\log(J+1)/C+\log 2)}{\log(J+1)}\right)\right)\geq\frac{q^{\prime}(q)-q}{2}\left(1-\frac{1}{C}-\frac{\log 2}{\log 3}\right)\geq\frac{q^{\prime}(q)-q}{8},$
for an appropriate value of the constant $C>0$.
We analyze $q^{\prime}(q)-q$ for $q^{\prime}(q)$ satisfying (14) as in Blanchard et al. (2024) and consider two regimes for $h(q\|q^{\prime})+h(q^{\prime}\|q)$. For any $0<q\leq q^{\prime}\leq\frac{1}{2}$, we have
$\frac{(q^{\prime}-q)^{2}}{q^{\prime}}\leq h(q\|q^{\prime})+h(q^{\prime}\|q)\leq\frac{2(q^{\prime}-q)^{2}}{q}.$ (15)
So, by the right inequality in (15),
$q^{\prime}(q)-q\geq\sqrt{\frac{q\left(h(q\|q^{\prime}(q))+h(q^{\prime}(q)\|q)\right)}{2}}=\sqrt{\frac{q\log(J+1)}{4Cn}}=\sqrt{\frac{S(p_{J+1})}{4Cn}}=\sqrt{\frac{s}{4Cn}}.$ (16)
In addition, by the left inequality in (15),
$q^{\prime}\leq q+\sqrt{\frac{q^{\prime}\log(J+1)}{2Cn}}\leq\sqrt{q^{\prime}}\left(\sqrt{q}+\sqrt{\frac{\log(J+1)}{2Cn}}\right),$
which implies
$q^{\prime}\leq\left(\sqrt{q}+\sqrt{\frac{\log(J+1)}{2Cn}}\right)^{2}$ $=q+2\sqrt{\frac{q\log(J+1)}{2Cn}}+\frac{\log(J+1)}{2Cn}$
$=q\left(1+2\sqrt{\frac{\log(J+1)}{2Cnq}}+\frac{\log(J+1)}{2Cnq}\right).$
Since by assumption $Cqn=\frac{Cn}{\frac{t}{s}\log(\frac{t}{s})}\geq c^{\prime}\log 2$, we have that for a sufficiently large constant $c^{\prime}$,
$q^{\prime}\leq\frac{q\log(J+1)}{\log 2}.$
This verifies (11) and establishes the term $c\sqrt{\frac{s}{n}}$ in (8).
Next, we assume $\frac{t}{n}\geq c\sqrt{\frac{s}{n}}$. For any $20\leq q\leq q^{\prime}\leq 1/$ 2 we have $h(q\|q^{\prime})\leq h(q^{\prime}\|q)$, and for $q^{\prime}\geq 9q$, it holds that (see, e.g., Blanchard et al. (2024))
$\frac{q}{2}\cdot\tfrac{q^{\prime}-q}{q}\log\tfrac{q^{\prime}-q}{q}\leq h(q\lVert q^{\prime})+h(q^{\prime}\lVert q)\leq 4q\cdot\tfrac{q^{\prime}-q}{q}\log\tfrac{q^{\prime}-q}{q}.$ (17)
For $z\geq\mathrm{e}$, the solution $x$ to the equation $x\log x=z$ satisfies $x\geq\frac{z}{\log z}$ (Corless et al., 1996). Taking $c>0$ sufficiently large such that
$z=\frac{\log(J+1)}{2Cqn}=\frac{t^{2}}{2Cns}\log^{2}\left(\tfrac{t}{s}\right)\geq\frac{c^{2}}{2C}\geq\mathrm{e},$
we have that $q^{\prime}(q)$ satisfies (14) if $q^{\prime}(q)-q\geq 8q$ and
$\frac{q^{\prime}(q)-q}{q}\geq\frac{\frac{\log(J+1)}{2Cqn}}{\log\frac{\log(J+1)}{2Cqn}};$
namely,
$q^{\prime}(q)-q$ $\geq\frac{\log(J+1)}{2Cn\log\left(\frac{\log(J+1)}{2Cqn}\right)}=\frac{T(p_{J+1})}{2Cn}\cdot\frac{\log\frac{1}{q}}{\log\frac{\log(J+1)}{2Cqn}}$
$=\frac{T(p_{J+1})}{2Cn}\cdot\frac{\log\frac{1}{q}}{\log\frac{s}{2Cnq^{2}}}\geq\frac{T(p_{J+1})}{4Cn}=\frac{t}{4Cn}\cdot\left(1+\frac{\log\frac{t}{s}}{\log\log\frac{t}{s}}\right)^{-1},$
where in the last inequality we used the fact that $\frac{s}{2Cn}\leq 1$. Lastly, we verify that $q^{\prime}(q)$ is such that (11) holds, namely, $\frac{\log\frac{1}{q}}{\log\frac{1}{q^{\prime}(q)}}\leq\frac{\log(J+1)}{\log 2}.$ The left inequality in (17) implies that $q^{\prime}(q)\leq q+C^{\prime}\frac{t}{n}$ for some constant $C^{\prime}$. Putting this and $q=\frac{1}{\frac{t}{s}\log\frac{t}{s}}$ and $\log(J+1)=t\log\frac{t}{s}$, we have that (11) holds if
$\frac{\log(\frac{t}{s}\log\frac{t}{s})}{\log(\frac{t}{s}\log\frac{t}{s})-\log(1+\frac{t^{2}}{Cns}\log\frac{t}{s})}\leq\frac{t\log\frac{t}{s}}{\log 2}.$
This is satisfied when
$n\geq\frac{\frac{t^{2}}{Cs}\log\frac{t}{s}}{\frac{t}{s}\log\frac{t}{s}\cdot\mathrm{e}^{-\frac{t\log\frac{t}{s}}{\log 2}}-1}.$
Finally, we consider the case where $t\geq n$. We repeat the arguments in the proof of Theorem 1 to show that in this case the minimax rate is bounded from below by a constant. Taking any $p^{*}\in\mathcal{P}_{s,t}$ such that $T(p^{*})\geq n$, and assuming without loss of generality that $p^{*}$ is non-increasing, let $j^{\prime}$ be such that
$T(p^{*})\geq\frac{\log(1+j^{\prime})}{\log(1/\dot{p}_{j^{\prime}}^{*})}.$
Then
$\sum_{j=1}^{\infty}(\dot{p}_{j}^{*})^{n}\geq\sum_{j=1}^{\infty}(\dot{p}_{j}^{*})^{T(p)}\geq\sum_{j=1}^{j^{\prime}}(\dot{p}_{j}^{*})^{\frac{\log(1+j^{\prime})}{\log(1/(\dot{p}_{j}^{*}))}}\geq j^{\prime}(\dot{p}_{j^{\prime}}^{*})^{\frac{\log(1+j^{\prime})}{\log(1/(\dot{p}_{j^{\prime}}^{*}))}}=\frac{j^{\prime}}{1+j^{\prime}}\geq\frac{1}{2}.$
As in the proof of Theorem 1, applying Lemma 2 with
$A_{j}=\left\{(X^{(1)}_{j},\dots,X^{(n)}_{j})=(1,1,\dots,1),\text{ but }1-\dot{p}_{j}^{*}\neq p_{j}^{(Y)}\right\},$
where $Y_{j}\sim\operatorname{Bernoulli}(1/2)$ and $p_{j}^{(Y_{j})}=Y_{j}\dot{p}_{j}^{*}+(1-Y_{j})(1-\dot{p}_{j}^{*})$, we get that the minimax rate is lower bounded by a universal constant. $\square$
3.3 Proof of Theorem 3
We aim to prove that the family
$\mathcal{P}:=\dot{\mathsf{LGC}}\cup\left\{(c,c,\dots):c\in[0,1]\right\}$
is learnable by an estimator $\tilde{p}_{n}$. Choose some $p\in\mathcal{P}$. The general strategy is to construct an estimator that can distinguish between cases where $T(p)=\infty$ and cases where $T(p)<\infty$, based on the sample.
Step 1: Testing if $T(p)=\infty$.
We begin by defining a test $\Phi$ to check whether $T(p)=\infty$. The idea is to check the first half of the sequence $X^{(1)}_{j},X^{(2)}_{j},\ldots,X^{(n)}_{j}$ are all ones and the second half are all zeros. Formally, the test function is defined as:
$\Phi(X^{(1)},X^{(2)},\ldots,X^{(n)})=\mathbf{1}\left(\limsup_{j\to\infty}B_{j}\right),$
where we define the event
$E_{j}:=\left\{X^{(i)}_{j}=0\text{ for }i\leq\frac{n}{2}\text{ and }X^{(i)}_{j}=1\text{ for }i>\frac{n}{2}\right\}.$
Note that, for each $j$, we have
$\mathbb{P}\left(E_{j}\right)=p_{j}^{{\left\lfloor n/2\right\rfloor}}\left(1-p_{j}\right)^{{\left\lceil n/2\right\rceil}},$
and because $\left\{E_{j}\right\}_{j}$ are independent, by the two Borell-Cantelli lemmas, $\Phi=1$ almost surely if and only if
$\sum_{j=1}^{\infty}p_{j}^{{\left\lfloor n/2\right\rfloor}}\left(1-p_{j}\right)^{{\left\lceil n/2\right\rceil}}=\infty.$
The above sum can be estimated by the following sums,
$\sum_{j=1}^{\infty}\dot{p}_{j}^{n}\leq\sum_{j=1}^{\infty}p_{j}^{{\left\lfloor n/2\right\rfloor}}\left(1-p_{j}\right)^{{\left\lceil n/2\right\rceil}}\leq\sum_{j=1}^{\infty}\dot{p}_{j}^{{\left\lfloor n/2\right\rfloor}}.$ (18)
Step 2: Consistency of the Test.
We now show that the test $\Phi$ is consistent. First, assume $T(p)=\infty$, which means $T(\dot{p}_{\downarrow 0})=\infty$, then by Cohen and Kontorovich (2023, Lemma 3), we have $\sum_{j=1}^{\infty}\dot{p}_{j}^{n}=\infty$ for all $n$, then by (18) we have $\Phi=1$ almost surely.
On the other hand, if $T(p)<\infty$, again by Cohen and Kontorovich (2023, Lemma 3), we have $\sum_{j=1}^{\infty}\dot{p}_{j}^{n}<\infty$ for large enough $n$, which means that for large enough $n$ we have $\Phi=0$ almost surely, as before.
Step 3: Defining the Estimator.
Once the test $\Phi$ has been applied, we define the estimator $\tilde{p}_{n}$ as follows:
$\tilde{p}_{n}(j)=\begin{cases}\frac{1}{n}\sum_{i=1}^{n}\hat{p}_{n}(j),&\text{if }\Phi(X_{1},\dots,X_{n})=1,\\ \hat{p}_{n}(j),&\text{otherwise}.\end{cases}$
In words, if the test $\Phi$ indicates that $T(p)=\infty$, we use the average of all $\hat{p}_{n}(j)$ (as this is consistent with the assumption that $p$ is a constant sequence). If $\Phi$ indicates that $T(p)<\infty$, we use the MLE $\hat{p}_{n}(j)$ directly.
Step 4: Consistency of the Estimator.
We now verify that the estimator $\tilde{p}_{n}$ is consistent.
If $p\in\mathsf{LGC}$, then $T(p)<\infty$, and the test $\Phi$ will eventually return 00. In this case, the estimator $\tilde{p}_{n}(j)$ is simply the MLE, which is known to be consistent for all $p\in\mathsf{LGC}$. Therefore, $\tilde{p}_{n}(j)\to p(j)$ as $n\to\infty$.
If $p=(c,c,\dots)$ for some constant $c\in[0,1]$, then $T(p)=\infty$, and the test $\Phi$ will always return $1$. In this case, the estimator $\tilde{p}_{n}(j)$ is the average of all $\hat{p}_{n}(j)$, which, by the law of large numbers, will converge to $c$. Thus, $\tilde{p}_{n}(j)\to c$ as $n\to\infty$.
Conclusion.
The estimator $\tilde{p}_{n}$ correctly learns all distributions in the family $\dot{\mathsf{LGC}}\cup\left\{(c,c,\dots):c\in[0,1]\right\}$, completing the proof.
$\square$
3.4 Auxiliary lemmas
Lemma 1 (Yu (1997)).
For $r\geq 2$, let $\nu_{1},\nu_{2},...,\nu_{r}$ be a collection of $r$ probability measures with some parameter of interest $\theta(\nu)$ taking values in pseudo-metric space $(\Theta,\rho)$ such that for all $j\neq k$,
$\rho(\theta(\nu_{j}),\theta(\nu_{k}))\geq\alpha$
and
$D_{\textup{KL}}(\nu_{j}\|\nu_{k})\leq\beta.$
Then
$\inf_{\hat{\theta}}\max_{k\in[r]}\mathop{\mathbb{E}}_{Z\sim\nu_{k}}\rho(\hat{\theta}(Z),\theta(\nu_{k}))\geq\frac{\alpha}{2}\left(1-\left(\frac{\beta+\log 2}{\log r}\right)\right),$
where the infimum is over all estimators $\hat{\theta}:Z\mapsto\Theta$.
Lemma 2 (Van Handel (2014) Problem 5.1a).
If $A_{1},\ldots,A_{N}$ are independent events, then
$(1-\mathrm{e}^{-1})\left[1\wedge\sum_{k=1}^{N}\mathbb{P}(A_{k})\right]$$\leq$ $\mathbb{P}\left(\bigcup_{k=1}^{N}A_{k}\right).$